OpenStudy (anonymous):

Brian took eight years to pay off his $71,900 loan. The loan had an interest rate of 8.16%, compounded quarterly. If Brian paid quarterly and made the same payment every time, how much was each payment that he made? a. $2,342.66 b. $3,081.54 c. $1,022.28 d. $1,466.76

OpenStudy (anonymous):

Ok we know that the formula for compound intrest is \[A=P(1+\frac{ r }{ n })^(nt)\] So P=$71,900 r=0.0816 n=4(because it is quarterly) and t=8 (because it took him 8 years to pay off his dept

OpenStudy (anonymous):

Sorry about the messed up equation but on the last part its sposed to be to the power of (nt)

OpenStudy (anonymous):

So all we have to do is plug in the values like so here \[A=(71,900)(1+\frac{ 0.0816 }{ 4 })^{6*4}\]

OpenStudy (anonymous):

Now calculate this and tell me what you get

OpenStudy (anonymous):

137,209.217

OpenStudy (anonymous):

Ok lemme check your answer

OpenStudy (anonymous):

I actually calculated $116740 Maybe take each part one by one \[(1+\frac{ .0816 }{ 4 })^{24}=1.6236\] then \[71900*1.6236=116740\]

OpenStudy (anonymous):

thats not any of my choices though

OpenStudy (anonymous):

Hmm how odd what are the choices

OpenStudy (anonymous):

a. $2,342.66 b. $3,081.54 c. $1,022.28 d. $1,466.76

OpenStudy (anonymous):

Ohhh your calculating each payment ok

OpenStudy (anonymous):

so 3months*8years=24months so 116740/24=4864

OpenStudy (anonymous):

Hmm :/

OpenStudy (anonymous):

makes no sense to me

OpenStudy (anonymous):

I dont know aht I did wrong since we calculated the total amount with intrest then I divided it by 4 for each payment

OpenStudy (anonymous):

quater

OpenStudy (anonymous):

Hold on I think i got it

OpenStudy (anonymous):

Drats I got to go. Sorry dude I dont know what to do

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

it fine